OpenStudy (anonymous):

Identify the 42nd term of an arithmetic sequence where a1 = -12 and a27 = 66. (2 points) 70 72 111 114

OpenStudy (welshfella):

common difference = (a27 - a1) / 26

OpenStudy (welshfella):

= (66 -(-12) / 26 = 78/26 = ?

OpenStudy (welshfella):

if common difference = d 42nd term = -12 + 41d

OpenStudy (welshfella):

- plug in the value for d you;ve just found

OpenStudy (anonymous):

is the common difference a 3?